OpenStudy (anonymous):

Which proportion has a solution of 8? a.thirty over twelve equals n over forty-five b.twelve over thirty equals n over forty-five c.twelve over n equals forty-five over thirty d.twelve over thirty equals forty-five over n

OpenStudy (ryan123345):

lets start with the first choice \[\frac{ 30 }{ 12 }=\frac{ n }{ 45 }\]

OpenStudy (ryan123345):

whats 45 x 30 ?

OpenStudy (anonymous):

1350

OpenStudy (ryan123345):

ok and 12 x n = 12n so you get 12n=1350 whats 1350 divided by 12?

OpenStudy (anonymous):

112.5

OpenStudy (ryan123345):

ok so A is not right

OpenStudy (ryan123345):

lets try B. \[\frac{ 12 }{ 30 }=\frac{ n }{ 45 }\]whats 12 x 45?

OpenStudy (anonymous):

540

OpenStudy (ryan123345):

now whats 540 divided by 30?

OpenStudy (anonymous):

18

OpenStudy (ryan123345):

ok thats wrong. Lets try C . \[\frac{ 12 }{ n }=\frac{ 45 }{ 30 }\]

OpenStudy (anonymous):

its c thx

OpenStudy (ryan123345):

360 divided by 45 is? c: good job :)